40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jan 31 20:06:22 2013, Evan, Update, Locking, PRM/PR2 cavity 
    Reply  Thu Jan 31 23:48:45 2013, Koji, Update, Locking, PRM/PR2 cavity 
       Reply  Fri Feb 1 09:33:11 2013, Jamie, Update, Locking, PRM/PR2 cavity 
          Reply  Fri Feb 1 15:12:53 2013, Jenne, Update, Locking, PRM/PR2 cavity 
             Reply  Fri Feb 1 19:55:33 2013, Jenne, Update, Locking, PRM/PR2 cavity 
                Reply  Fri Feb 1 23:12:42 2013, Koji, Update, Locking, PRM/PR2 cavity 
                Reply  Fri Feb 1 23:52:59 2013, rana, Update, Locking, PRM/PR2 cavity 
Message ID: 7980     Entry time: Thu Jan 31 23:48:45 2013     In reply to: 7978     Reply to this: 7981
Author: Koji 
Type: Update 
Category: Locking 
Subject: PRM/PR2 cavity 

Wow! What's happened?

As the video showed good quality of resonances, I stopped by at the 40m on the way back home.

I looked at the error signals and found that they indicate high finesse and clear resonance of the sidebands.
The lock was immediate once the gain is set to be -0.004 (previous 0.05ish). This implies the optical gain is ~10 times larger than the previous configration.
The alignment was not easy as POPDC was saturated at ~27000I leave this as a daytime job.

As I misaligned the PRM, I could see that the lock hopped into the next higher order. i.e .from TEM00 to TEM01, from TEM01 to TEM02, etc
This means that the modes are closely located each other, but sufficiently separated to sustain each mode.

I definitely certify that cavity scans will give us meaningful information about the cavity.

ELOG V3.1.3-